OpenStudy (anonymous):

cos^-1 (3/8) or arccos 3/8

OpenStudy (acal21):

they are the same thing

OpenStudy (anonymous):

any answers u got

OpenStudy (anonymous):

yes i know they are the same thing, some people only reconize one way. i understand how to input into a calculator, but the teacher ask us to show work.

OpenStudy (anonymous):

i know the answer is 1.19, but i need to show work on paper

OpenStudy (anonymous):

@acal21 , you only came to say its the same problem? i assume you dont know how to show work for it either!

OpenStudy (anonymous):

@ozzy22 , can you show me how to work out the problem with no calculator?

OpenStudy (anonymous):

this is actually an angle difference formula for tan problem tan (A - B) = (tan A - tan B) / (1 + tan A tan B) A = cos^-1 (-3/5) ==> x = -3 and r = 5; therefore, y = 4 (from x^2 + y^2 = r^2) tan A = -4/3 B = arcsin (8/17) ==> y = 8, r = 17, and x = 15 (again, from x^2 + y^2 = r^2) tan B = 8/15 plugging into formula: tan (A - B) = (-4/3 - 8/15) / (1 + (-4/3)(8/15)) tan (A - B) = (-20/15 - 8/15) / (1 - 32/45) tan (A - B) = (-28/15) / (13/45) = (-28/15) * (45/13) = (-3*28) / 13 = -84/13 tan (A - B) = -84/13

OpenStudy (anonymous):

@ozzy22 so what do you have as the final answer?

OpenStudy (anonymous):

-84/13?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the correct answr for this problem is 1.19

OpenStudy (anonymous):

how did u get that answer POST

OpenStudy (anonymous):

my book has the answers to the odd problems, it only shows the key strokes for the solution, do you want that?

OpenStudy (anonymous):

yes pls

OpenStudy (anonymous):

one moment, have to take pic with phone,email to laptop then upload here

OpenStudy (anonymous):

K NP

OpenStudy (anonymous):

@ozzy22

OpenStudy (anonymous):

1.20 rounded in two places

OpenStudy (anonymous):

idk man, im completely lost on this problem

OpenStudy (anonymous):

oh round to 2 decimal places in instructions

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry about the first answer

OpenStudy (anonymous):

no problem, im not dwelling on mistakes. just looking for a way to show work for this on paper

OpenStudy (anonymous):

\[3\div8=1.199 \]

OpenStudy (anonymous):

with cos-1

OpenStudy (anonymous):

then round .199 witch